North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size

Wednesday 4th October 2017

Bit of a late start, about 10am, under a cloudy sky with a hint of dampness. Bill Dobson, Ian Pearson, Nigel Hall, and myself were present, along with Jon Bradley making a welcome reappearance after his visit to the USA in September. First job was getting some water so we could all have our usual morning tea, as the bottles had not been filled up by the team at the end of Saturday.

As Ian reported earlier in the week, the Q6 has now been shunted on to the wheel drops, and the leading set of wheels were fitted by Shaun Bowler and his MPD team on Tuesday. Work was still ongoing yesterday with springs and other items being fitted . While this work was being done, Jon and Ian removed the steam brake piston and cover from the cab of the locomotive and took it to Deviation Shed in the wheelbarrow for cleaning and inspection. It was found that the piston shaft at the rear end had some burrs on it which could have caused its reluctance to come apart during previous attempts to remove it from the cylinder. These burrs were ground out along with the shoulder part of the rod. Jon cleaned the cover and found a rough part inside the brass bush which could have also contributed to the piston rod being stubborn to move. This was filed smooth and a trial fit of the cover over the piston shaft proved to be a good fit. While Bill was supervising and delegating jobs, and checking work at the wheel drops with Shaun and his team, I was given the job of freeing the L/H tender water valve linkage which was found to be seized on Saturday. With the help of Ian who applied some heat to the seized crank (eat your heart out Trevor!), it was freed up, and by the end of the day, after it had cooled, the linkages were reconnected and found to be working smoothly.

Nigel had to go to Whitby to buy some Isopon filler before he could start prepping the cylinder cladding sheets. In spite of this, by the end of the day he had successfully cleaned, filled, rubbed down and primed the left and right cylinder cladding sheets. Bill took off the L/H tender axle box cover to examine the leakage from this box. He was unable to remove the oil pad fully so the pad was put back in pending further investigation. Ian removed the 3 pins from the steam brake piston linkage arm and got them shot blasted. As the NDT man was on site doing jobs for the MPD we got him to come and NDT our pins. These were all found to be OK and refitted to the brake linkage.

Lunch was had in Deviation Shed today, with not many visitors in spite of Repton being parked outside all day out of steam. One commented that, in spite of being a regular visitor to the Railway, he had not realised that you could come up this far - how are you getting on with the new signage Bryan? In spite of this, the bell on the donation box was ringing merrily throughout the afternoon.

Later in the afternoon, at the wheel drops, the trailing wheel set was being lifted up into the horn guides when the hydraulics failed and the job was stopped . The wheel drops are now non operational and awaiting repairs.

The only other excitement of the day was the delivery of the new cylinder block casting for No 29 (along with the original and the patterns). It is now in the MPD awaiting assessment by the company who are in line to do the machining. I should also report though that those strange foreign machines from the West have left the Railway, and things are back to normal motive power wise!